# Navigate Beyond Shortcuts: Debiased Learning through the Lens of Neural Collapse

### v1

@@ 1 Introduction
To achieve efficient model debiasing, we follow the inspiration of prime training, and encourage the model to skip the active learning of shortcut correlations. The primes are often provided as additional supervisory signals to redirect the model’s reliance on shortcuts, which helps improve generalization in image classification and CARLA autonomous driving [33]. To rectify models’ attention on the intrinsic correlations, we define the primes with a training-free simplex ETF structure, which approximates the “optimal" shortcut features and guides the model to pursue unbiased classification from the beginning of training. Our method is free of auxiliary models or additional optimization of prime features. Experimental results also substantiate its state-of-the-art debiasing performance on both synthetic and real-world biased datasets.
@@ 2 Related Works
Avoid-shortcut Learning. The recent inspiration of avoid-shortcut learning aims to postpone, or even prevent the learning of shortcut relations in model training. With well-crafted contrastive samples [27, 29, 28] or artificial shortcut signals [33, 42], avoid-shortcut learning has demonstrated its efficacy in image classification, autonomous driving and question answering models. One of the representative methods is named prime training, which provides richer supervisory signals of key input features (i.e., primes) to guide the establishment of correct correlations, therefore improving generalization on OOD samples [33]. In this work, we leverage the approximated “optimal" shortcuts as primes to encourage the models to bypass shortcut learning.
@@ 3.1 Neural Collapse Phenomenon
where $\mathbf{M}=[\mathbf{m}_{1},...,\mathbf{m}_{K}]\in\mathbb{R}^{d\times K}$, and $\mathbf{P}\in\mathbb{R}^{d\times K}$ is an orthogonal matrix which satisfies $\mathbf{P}^{\mathrm{T}}\mathbf{P}=\mathbf{I}_{K}$, with $\mathbf{I}_{K}$ denotes the identity matrix and $\mathbf{1}_{K}$ denotes the all-ones vector. Within the ETF structure, all vectors have the maximal pair-wise angle of $-\frac{1}{K-1}$, namely the maximal equiangular separation.
@@
NC3: Self duality. The classifier weights $\mathbf{w}_{k}$ will align with the corresponding normalized class means $\mathbf{{\tilde{z}}}_{k}$, which satisfies $\mathbf{{\tilde{z}}}_{k}=\mathbf{w}_{k}/||\mathbf{w}_{k}||$.
NC4: Simplification to nearest class center. After convergence, the model’s prediction will collapse to simply choosing the nearest class mean to the input feature (in standard Euclidean distance). The prediction of $\mathbf{z}$ could be denoted as $\arg\max_{k}\langle\mathbf{z},\mathbf{w}_{k}\rangle=\arg\min_{k}||\mathbf{z}-\mathbf{\overline{z}}_{k}||$.
@@ 3.2 Neural Collapse Observation on Biased Dataset
Metrics Computational details NC1 ${\rm\mathcal{NC}_{1}}=\frac{1}{K}{\rm Tr}(\Sigma_{W}\Sigma_{B}^{\dagger})$, where ${\rm Tr}$ is the trace of matrix and $\Sigma_{B}^{\dagger}$ denotes the pseudo-inverse of $\Sigma_{B}$ $\Sigma_{B}=\frac{1}{K}\sum\limits_{k\in[K]}(\mathbf{\overline{z}}_{k}-\mathbf{z}_{G})(\mathbf{\overline{z}}_{k}-\mathbf{z}_{G})^{\mathrm{T}}$, $\Sigma_{W}=\frac{1}{K}\sum\limits_{k\in[K]}\frac{1}{n_{k}}\sum\limits_{i=1}^{n_{k}}(\mathbf{z}_{k,i}-\mathbf{\overline{z}}_{k})(\mathbf{z}_{k,i}-\mathbf{\overline{z}}_{k})^{\mathrm{T}}$ NC2 ${\rm\mathcal{NC}_{2}}=\Big{\lVert}\frac{\mathbf{W}\mathbf{W}^{\mathrm{T}}}{||\mathbf{W}\mathbf{W}^{\mathrm{T}}||_{\mathcal{F}}}-\frac{1}{\sqrt{K-1}}(\mathbf{I}_{K}-\frac{1}{K}\mathbf{1}_{K}\mathbf{1}_{K}^{\mathrm{T}})\Big{\rVert}_{\mathcal{F}}$ NC3 ${\rm\mathcal{NC}_{3}}=\Big{\lVert}\frac{\mathbf{W}\mathbf{\overline{Z}}}{||\mathbf{W}\mathbf{\overline{Z}}||_{\mathcal{F}}}-\frac{1}{\sqrt{K-1}}(\mathbf{I}_{K}-\frac{1}{K}\mathbf{1}_{K}\mathbf{1}_{K}^{\mathrm{T}})\Big{\rVert}_{\mathcal{F}}$, where $\mathbf{\overline{Z}}=[\mathbf{\overline{z}}_{1}-\mathbf{z}_{G},...,\mathbf{\overline{z}}_{K}-\mathbf{z}_{G}]$
@@ 4.1 Motivation
Following the previous analysis, we highlight the importance of redirecting the model’s emphasis from simple shortcuts to intrinsic relations. Since the models can be easily misled by shortcuts in the training distribution, is it feasible to supply a “perfectly learned" shortcut feature, to deceive the models into skipping the active learning of shortcuts, and directly focusing on the intrinsic correlations?
We observe in Fig. 2 that the NC1-NC3 metrics show a rapid decrease during shortcut learning, but remain stable in the subsequent training epochs. However, if the training distribution does follow the shortcut correlation (with no obstacle from bias-conflicting samples), the convergence will end up with the optimal structure of simplex ETF, just as the results on unbiased datasets. This inspires us to approximate the “perfectly learned" shortcut features with a simplex ETF structure, which requires no additional training and represents the optimal geometry of feature space.
Therefore, following the outstanding performance of prime training in OOD generalization [33], we introduce the approximated “perfect" shortcuts as the primes for debiased learning. The provided shortcut primes are constructed with a training-free simplex ETF structure, which encourages the models to directly capture the intrinsic correlations, therefore exhibit superior generalizability and convergence properties in our experiments.
@@ 4.2 Avoid-shortcut Learning with Neural Collapse
Building upon our motivation of avoid-shortcut learning, the illustration of the proposed ETF-Debias is shown in Fig. 3. The debiased learning framework can be divided into three stages: prime construction, prime training, and unbiased classification. Firstly, a prime ETF will be constructed to approximate the “perfect" shortcut features. Then during the prime training, the model will be guided to directly capture the intrinsic correlations with the prime training and the prime reinforcement regularization. In evaluation, we rely on the intrinsic correlations to perform unbiased classification. The details are as follows.
Prime construction. When constructing the prime ETF, we first randomly initialize a simplex ETF as $\mathbf{M}\in\mathbb{R}^{d\times B}$, which satisfies the definition in Eq. 1. The dimension $d$ is the same as the learnable features, and the number of vectors in $\mathbf{M}$ is determined by the categories of bias attributes $\mathbf{b}_{i}\in\{1,...,B\}$, which are pre-defined in the training distribution. After initialization, the vertices of prime ETF $[\mathbf{m}_{1},...,\mathbf{m}_{B}]$ are considered as the approximation of the “perfect" shortcut features for each attribute, which serve as the prime features for avoid-shortcut training. During training, the prime features will be retrieved based on the bias attribute $b$ of each input sample.
@@
In essence, we provide a pre-defined prime feature for each training sample based on its bias attribute. The prime features, with a strong correlation with the bias attributes, can be viewed as the optimal solution to shortcut learning. By leveraging the already “perfect" representation of shortcut correlations, the model will be forced to explore the intrinsic correlations within the training distribution. The prime-guided mechanism targets at the fundamental issue of biased classification, without inducing extra training costs.
Prime reinforcement regularization. Given the prime features, the model is encouraged to grasp the intrinsic correlation of the training distributions. However, we raise another potential risk that, despite the provided “perfectly learned" shortcut features, the model may still pursue the easy-to-follow shortcuts, leading to the redundancy between the learnable feature $\mathbf{z}_{i,b}$ and the fixed $\mathbf{m}_{b}$. We point out that the model may not establish a strong correlation between the prime features and the bias attributes, and continues to optimize the learnable features for the missing connections.
@@ 5 Experiments
Dataset Ratio(%) Vanilla LfF[23] LfF+BE[19] EnD[30] SD[42] DisEnt[18] Selecmix[13] ETF-Debias Colored MNIST 0.5 32.22$\pm$0.13 57.78$\pm$0.81 69.69$\pm$1.99 35.93$\pm$0.40 56.96$\pm$0.37 68.83$\pm$1.62 70.53$\pm$0.46 71.63$\pm$0.28 (+1.10) 1.0 48.45$\pm$0.06 72.29$\pm$1.69 80.90$\pm$1.40 49.32$\pm$0.58 72.46$\pm$0.18 79.49$\pm$1.44 83.34$\pm$0.37 81.97$\pm$0.26 (-1.37) 2.0 58.90$\pm$0.12 79.51$\pm$1.82 84.90$\pm$1.14 65.58$\pm$0.46 79.37$\pm$0.46 84.56$\pm$1.19 85.90$\pm$0.23 86.00$\pm$0.03 (+0.10) 5.0 74.19$\pm$0.04 83.96$\pm$1.44 90.28$\pm$0.18 80.70$\pm$0.17 88.89$\pm$0.21 88.83$\pm$0.15 91.27$\pm$0.31 91.36$\pm$0.21 (+0.09) Corrupted CIFAR-10 0.5 17.06$\pm$0.12 31.00$\pm$2.67 23.68$\pm$0.50 14.30$\pm$0.10 36.66$\pm$0.74 30.12$\pm$1.60 33.30$\pm$0.26 40.06$\pm$0.03 (+3.40) 1.0 21.48$\pm$0.55 34.33$\pm$1.76 30.72$\pm$0.12 20.17$\pm$0.19 45.66$\pm$1.05 35.28$\pm$1.39 38.72$\pm$0.27 47.52$\pm$0.26 (+1.86) 2.0 27.15$\pm$0.46 39.68$\pm$1.15 42.22$\pm$0.60 30.10$\pm$0.54 50.11$\pm$0.69 40.34$\pm$1.41 47.09$\pm$0.17 54.64$\pm$0.42 (+4.53) 5.0 39.46$\pm$0.58 53.04$\pm$0.76 57.93$\pm$0.58 45.85$\pm$0.21 62.43$\pm$0.57 49.99$\pm$0.84 54.69$\pm$0.29 65.34$\pm$0.60 (+2.91)
@@
Dataset Ratio(%) Vanilla LfF[23] LfF+BE[19] EnD[30] SD[42] DisEnt[18] Selecmix[13] ETF-Debias Biased FFHQ 0.5 53.27$\pm$0.61 65.60$\pm$2.27 67.07$\pm$2.37 55.93$\pm$1.62 65.60$\pm$0.20 63.07$\pm$1.14 65.00$\pm$0.82 73.60$\pm$1.22 (+6.53) 1.0 57.13$\pm$0.64 72.33$\pm$2.19 73.53$\pm$1.62 61.13$\pm$0.50 69.20$\pm$0.20 68.53$\pm$2.32 67.50$\pm$0.30 76.53$\pm$1.10 (+3.00) 2.0 67.67$\pm$0.81 74.80$\pm$2.03 80.20$\pm$2.78 66.87$\pm$0.64 78.40$\pm$0.20 72.00$\pm$2.51 69.80$\pm$0.87 85.20$\pm$0.61 (+5.00) 5.0 78.87$\pm$0.83 80.27$\pm$2.02 87.40$\pm$2.00 80.87$\pm$0.42 84.80$\pm$0.20 80.60$\pm$0.53 83.47$\pm$0.61 94.00$\pm$0.72 (+6.60) Dogs & Cats 1.0 51.96$\pm$0.90 71.17$\pm$5.24 78.87$\pm$2.40 51.91$\pm$0.24 78.13$\pm$1.06 65.13$\pm$2.07 54.19$\pm$1.61 80.07$\pm$0.90 (+1.20) 5.0 76.59$\pm$1.27 85.83$\pm$1.62 88.60$\pm$1.21 79.07$\pm$0.28 89.12$\pm$0.18 82.47$\pm$2.86 81.50$\pm$1.06 92.18$\pm$0.62 (+3.06) BAR 1.0 68.00$\pm$0.43 68.30$\pm$0.97 71.70$\pm$1.33 68.25$\pm$0.19 67.33$\pm$0.35 69.30$\pm$1.27 69.83$\pm$1.02 72.79$\pm$0.21 (+1.09) 5.0 79.34$\pm$0.19 80.25$\pm$1.27 82.00$\pm$1.24 78.86$\pm$0.36 79.10$\pm$0.42 81.19$\pm$0.70 78.79$\pm$0.52 83.66$\pm$0.21 (+1.66)
@@ 5.3 Ablation Study
Ablation on the influence of ETF prime features. As illustrated before, we choose the vertices of ETF as the “perfectly learned" shortcut features, thus redirecting the model’s attention to intrinsic correlations. To demonstrate the efficacy of ETF prime features, we compare the results of randomly initialized prime features with the same dimension as the ETF-based ones. As shown in Fig. 4(b), the randomly initialized primes suffer a severe performance degradation, underscoring the advantages of ETF-based prime features in approximating the optimal structure.
@@ A.2 Analysis of ETF-Debias
When trained on biased datasets, we assume each class $[1,...,K]$ is strongly correlated with a bias attribute $[b_{1},...,b_{K}]$. Based on the mechanism of prime training, we denote the $i$-th feature of the $k$-th class as $\widetilde{\mathbf{z}}_{k,i}=[\mathbf{z}_{k,i},\mathbf{m}_{i,b}]\in\mathbb{R}^{2\times d}$, where $\mathbf{z}_{k,i}\in\mathbb{R}^{d}$ represents the learnable features, and $\mathbf{m}_{i,b}\in\mathbb{R}^{d}$ represents the prime features retrieved based on the bias attribute $b$ of the input sample. With the definition, a bias-aligned sample of the $k$-th class $\mathbf{x}_{k,i}$ will have its feature in form of $\widetilde{\mathbf{z}}_{k,i}=[\mathbf{z}_{k,i},\mathbf{m}_{b_{k}}]$, and a bias-conflicting sample will have its feature as $\widetilde{\mathbf{z}}_{k,i}=[\mathbf{z}_{k,i},\mathbf{m}_{b_{k^{\prime}}}],k^{\prime}\neq k$. To keep the same form, we denote the classifier weight as $\widetilde{\mathbf{w}}_{k}=[\mathbf{w}_{k},\mathbf{a}_{k}]\in\mathbb{R}^{2\times d}$, where $\mathbf{w}_{k}\in\mathbb{R}^{d}$ represents the weight for intrinsic correlations and $\mathbf{a}_{k}\in\mathbb{R}^{d}$ is the weight for shortcut features. In the detailed convergence result of Neural Collapse (Section C), we observe that, due to the fixed prime features and their strong correlation with the bias attributes, $\mathbf{a}_{k}$ will quickly collapse into its bias-correlated prime feature $\mathbf{m}_{b_{k}}$, and can be viewed as constant after just a few steps of training. Thus the cross-entropy loss can be re-written as:
@@
The predicted probabilities both satisfy 0<($10<(1-p_{k}(\widetilde{\mathbf{z}}_{k,i}))<$ 1 and $10<p_{k}(\widetilde{\mathbf{z}}_{k^{\prime},j})<$ 1, and the scaling step from Eq. A.6 to Eq. A.7 is based on the Jensen inequality. The terms $p_{k}^{(l)}$ and $p_{k}^{(b)}$ are respectively the predicted probabilities for class labels and bias attributes, calculated with softmax:
@@ Appendix D Detailed Results of BAR dataset
Action Climbing Diving Fishing Vaulting Racing Throwing Average Vanilla 75.46$\pm$1.68 52.59$\pm$4.51 75.00$\pm$2.78 71.15$\pm$3.40 85.47$\pm$0.85 48.36$\pm$0.81 68.00$\pm$0.43 LfF [23] 72.52$\pm$8.73 67.16$\pm$8.96 69.44$\pm$5.56 71.14$\pm$3.18 79.77$\pm$1.98 40.84$\pm$5.08 68.30$\pm$0.97 LfF+BE [19] 82.78$\pm$3.86 62.96$\pm$5.88 62.96$\pm$4.24 82.07$\pm$6.79 82.62$\pm$3.85 43.42$\pm$2.27 71.70$\pm$1.33 EnD [30] 79.85$\pm$2.29 50.62$\pm$3.08 73.15$\pm$3.21 74.51$\pm$0.49 82.05$\pm$0.00 49.29$\pm$4.88 68.25$\pm$0.19 SD [42] 79.49$\pm$2.29 58.12$\pm$1.96 71.29$\pm$3.21 80.67$\pm$2.22 82.05$\pm$0.86 34.74$\pm$2.15 67.73$\pm$0.35 DisEnt [18] 79.85$\pm$4.57 54.56$\pm$0.85 75.00$\pm$5.56 73.67$\pm$4.15 87.17$\pm$2.57 44.13$\pm$2.15 69.30$\pm$1.27 Selecmix [13] 69.60$\pm$1.68 59.26$\pm$9.46 79.63$\pm$4.24 66.95$\pm$5.47 87.18$\pm$2.96 56.34$\pm$5.64 69.83$\pm$1.02 ETF-Debias 83.15$\pm$1.27 63.70$\pm$4.86 76.85$\pm$4.25 74.79$\pm$5.11 87.46$\pm$1.98 50.23$\pm$4.53 72.79$\pm$0.21 (+1.09)
@@
Action Climbing Diving Fishing Vaulting Racing Throwing Average Vanilla 86.44$\pm$0.64 85.19$\pm$0.75 79.63$\pm$3.20 82.63$\pm$1.75 89.85$\pm$0.19 52.11$\pm$6.13 79.34$\pm$0.19 LfF [23] 84.24$\pm$2.54 84.69$\pm$4.93 75.92$\pm$8.93 80.67$\pm$2.22 89.74$\pm$0.86 52.58$\pm$6.50 80.25$\pm$1.27 LfF+BE [19] 86.77$\pm$2.92 85.96$\pm$1.91 78.30$\pm$5.33 82.11$\pm$5.76 90.16$\pm$1.29 55.70$\pm$1.95 82.00$\pm$1.24 EnD [30] 85.71$\pm$1.10 86.42$\pm$1.13 84.26$\pm$1.61 78.15$\pm$0.84 92.59$\pm$1.98 46.01$\pm$0.81 78.86$\pm$0.36 SD [42] 84.25$\pm$0.64 82.91$\pm$4.12 82.41$\pm$1.60 86.27$\pm$1.28 88.03$\pm$0.86 50.70$\pm$1.41 79.10$\pm$0.42 Selecmix [13] 84.98$\pm$5.08 79.51$\pm$2.38 81.48$\pm$5.78 85.43$\pm$0.97 92.02$\pm$0.50 49.30$\pm$3.73 78.79$\pm$0.52 DisEnt [18] 84.98$\pm$1.68 89.13$\pm$5.66 77.77$\pm$0.00 80.39$\pm$2.95 90.59$\pm$1.71 48.82$\pm$1.63 81.19$\pm$0.70 ETF-Debias 90.48$\pm$1.68 87.65$\pm$6.39 85.18$\pm$1.61 93.28$\pm$2.22 92.31$\pm$2.26 53.05$\pm$0.81 83.66$\pm$0.21 (+1.66)

### v2

@@ 1 Introduction
To achieve efficient model debiasing, we follow the inspiration of prime training, and encourage the model to skip the active learning of shortcut correlations. The primes are often provided as additional supervisory signals to redirect the model’s reliance on shortcuts, which helps improve generalization in image classification and CARLA autonomous driving [33]. To rectify models’ attention on the intrinsic correlations, we define the primes with a training-free simplex ETF structure, which approximates the “optimal” shortcut features and guides the model to pursue unbiased classification from the beginning of training. Our method is free of auxiliary models or additional optimization of prime features. Experimental results also substantiate its state-of-the-art debiasing performance on both synthetic and real-world biased datasets.
@@ 2 Related Works
Avoid-shortcut Learning. The recent inspiration of avoid-shortcut learning aims to postpone, or even prevent the learning of shortcut relations in model training. With well-crafted contrastive samples [27, 29, 28] or artificial shortcut signals [33, 42], avoid-shortcut learning has demonstrated its efficacy in image classification, autonomous driving and question answering models. One of the representative methods is named prime training, which provides richer supervisory signals of key input features (i.e., primes) to guide the establishment of correct correlations, therefore improving generalization on OOD samples [33]. In this work, we leverage the approximated “optimal” shortcuts as primes to encourage the models to bypass shortcut learning.
@@ 3.1 Neural Collapse Phenomenon
where $\mathbf{M}=[\mathbf{m}_{1},...,\mathbf{m}_{K}]\in\mathbb{R}^{d\times K}$, and $\mathbf{P}\in\mathbb{R}^{d\times K}$ is an orthogonal matrix which satisfies $\mathbf{P}^{\mathrm{T}}\mathbf{P}=\mathbf{I}_{K}$, with $\mathbf{I}_{K}$ denotes the identity matrix and $\mathbf{1}_{K}$ denotes the all-ones vector. Within the ETF structure, all vectors have the maximal pair-wise angle of $\frac{1}{K-1}$, namely the maximal equiangular separation.
@@
[…]
@@ 3.2 Neural Collapse Observation on Biased Dataset
[…]
@@ 4.1 Motivation
Following the previous analysis, we highlight the importance of redirecting the model’s emphasis from simple shortcuts to intrinsic relations. Since the models can be easily misled by shortcuts in the training distribution, is it feasible to supply a “perfectly learned” shortcut feature, to deceive the models into skipping the active learning of shortcuts, and directly focusing on the intrinsic correlations?
We observe in Fig. 2 that the NC1-NC3 metrics show a rapid decrease during shortcut learning, but remain stable in the subsequent training epochs. However, if the training distribution does follow the shortcut correlation (with no obstacle from bias-conflicting samples), the convergence will end up with the optimal structure of simplex ETF, just as the results on unbiased datasets. This inspires us to approximate the “perfectly learned” shortcut features with a simplex ETF structure, which requires no additional training and represents the optimal geometry of feature space.
Therefore, following the outstanding performance of prime training in OOD generalization [33], we introduce the approximated “perfect” shortcuts as the primes for debiased learning. The provided shortcut primes are constructed with a training-free simplex ETF structure, which encourages the models to directly capture the intrinsic correlations, therefore exhibit superior generalizability and convergence properties in our experiments.
@@ 4.2 Avoid-shortcut Learning with Neural Collapse
Building upon our motivation of avoid-shortcut learning, the illustration of the proposed ETF-Debias is shown in Fig. 3. The debiased learning framework can be divided into three stages: prime construction, prime training, and unbiased classification. Firstly, a prime ETF will be constructed to approximate the “perfect” shortcut features. Then during the prime training, the model will be guided to directly capture the intrinsic correlations with the prime training and the prime reinforcement regularization. In evaluation, we rely on the intrinsic correlations to perform unbiased classification. The details are as follows.
Prime construction. When constructing the prime ETF, we first randomly initialize a simplex ETF as $\mathbf{M}\in\mathbb{R}^{d\times B}$, which satisfies the definition in Eq. 1. The dimension $d$ is the same as the learnable features, and the number of vectors in $\mathbf{M}$ is determined by the categories of bias attributes $\mathbf{b}_{i}\in\{1,...,B\}$, which are pre-defined in the training distribution. After initialization, the vertices of prime ETF $\mathbf{m}_{1},...,\mathbf{m}_{B}]$ are considered as the approximation of the “perfect” shortcut features for each attribute, which serve as the prime features for avoid-shortcut training. During training, the prime features will be retrieved based on the bias attribute $b$ of each input sample.
@@
In essence, we provide a pre-defined prime feature for each training sample based on its bias attribute. The prime features, with a strong correlation with the bias attributes, can be viewed as the optimal solution to shortcut learning. By leveraging the already “perfect” representation of shortcut correlations, the model will be forced to explore the intrinsic correlations within the training distribution. The prime-guided mechanism targets at the fundamental issue of biased classification, without inducing extra training costs.
Prime reinforcement regularization. Given the prime features, the model is encouraged to grasp the intrinsic correlation of the training distributions. However, we raise another potential risk that, despite the provided “perfectly learned” shortcut features, the model may still pursue the easy-to-follow shortcuts, leading to the redundancy between the learnable feature $\mathbf{z}_{i,b}$ and the fixed $\mathbf{m}_{b}$. We point out that the model may not establish a strong correlation between the prime features and the bias attributes, and continues to optimize the learnable features for the missing connections.
@@ 5 Experiments
[…]
@@
[…]
@@ 5.3 Ablation Study
Ablation on the influence of ETF prime features. As illustrated before, we choose the vertices of ETF as the “perfectly learned” shortcut features, thus redirecting the model’s attention to intrinsic correlations. To demonstrate the efficacy of ETF prime features, we compare the results of randomly initialized prime features with the same dimension as the ETF-based ones. As shown in Fig. 4(b), the randomly initialized primes suffer a severe performance degradation, underscoring the advantages of ETF-based prime features in approximating the optimal structure.
@@ A.2 Analysis of ETF-Debias
When trained on biased datasets, we assume each class [1,,K][1,...,K] is strongly correlated with a bias attribute $b_{1},...,b_{K}]$. Based on the mechanism of prime training, we denote the $i$-th feature of the $k$-th class as $\widetilde{\mathbf{z}}_{k,i}=[\mathbf{z}_{k,i},\mathbf{m}_{i,b}]\in\mathbb{R}^{2\times d}$, where $\mathbf{z}_{k,i}\in\mathbb{R}^{d}$ represents the learnable features, and $\mathbf{m}_{i,b}\in\mathbb{R}^{d}$ represents the prime features retrieved based on the bias attribute $b$ of the input sample. With the definition, a bias-aligned sample of the $k$-th class $\mathbf{x}_{k,i}$ will have its feature in form of $\widetilde{\mathbf{z}}_{k,i}=[\mathbf{z}_{k,i},\mathbf{m}_{b_{k}}]$, and a bias-conflicting sample will have its feature as $\widetilde{\mathbf{z}}_{k,i}=[\mathbf{z}_{k,i},\mathbf{m}_{b_{k^{\prime}}}],k^{\prime}\neq k$. To keep the same form, we denote the classifier weight as $\widetilde{\mathbf{w}}_{k}=[\mathbf{w}_{k},\mathbf{a}_{k}]\in\mathbb{R}^{2\times d}$, where $\mathbf{w}_{k}\in\mathbb{R}^{d}$ represents the weight for intrinsic correlations and $\mathbf{a}_{k}\in\mathbb{R}^{d}$ is the weight for shortcut features. In the detailed convergence result of Neural Collapse (Section C), we observe that, due to the fixed prime features and their strong correlation with the bias attributes, $\mathbf{a}_{k}$ will quickly collapse into its bias-correlated prime feature $\mathbf{m}_{b_{k}}$, and can be viewed as constant after just a few steps of training. Thus the cross-entropy loss can be re-written as:
@@
The predicted probabilities both satisfy $0<(1-p_{k}(\widetilde{\mathbf{z}}_{k,i}))<1$ and $0<p_{k}(\widetilde{\mathbf{z}}_{k^{\prime},j})<1$, and the scaling step from Eq. A.6 to Eq. A.7 is based on the Jensen inequality. The terms $p_{k}^{(l)}$ and $p_{k}^{(b)}$ are respectively the predicted probabilities for class labels and bias attributes, calculated with softmax:
@@ Appendix D Detailed Results of BAR dataset
[…]
@@
[…]